### v1

# RAMAC: Multimodal Risk-Aware Offline Reinforcement Learning and the Role of Behavior Regularization

Abstract
In safety-critical domains where online data collection is infeasible, offline reinforcement learning (RL) offers an attractive alternative but only if policies deliver high returns without incurring catastrophic lower-tail risk. Prior work on risk-averse offline RL achieves safety at the cost of value conservatism and restricted policy classes, whereas expressive policies are only used in risk-neutral settings. Here, we address this gap by introducing the Risk-Aware Multimodal Actor-Critic (RAMAC) framework, which couples an expressive generative actor with a distributional critic. The RAMAC differentiates composite objective combining distributional risk and BC loss through the generative path, achieving risk-sensitive learning in complex multimodal scenarios. We instantiate RAMAC with diffusion and flow-matching actors and observe consistent gains in $\mathrm{CVaR}_{0.1}$ while maintaining strong returns on most Stochastic-D4RL tasks. Code: https://github.com/KaiFukazawa/RAMAC.git
1 Introduction
In high-stakes applications such as autonomous driving, robotics, finance, and healthcare, where real-life explorations may lead to catastrophic consequences, offline RL offers a safe approach for generating policies that not only maximize long-horizon returns but also tightly control risk (Levine et al., 2020). Recent expressive generative policies (Wang et al., 2023; Park et al., 2025; Koirala & Fleming, 2025) can capture multimodal behavior and thus excel in achieving high expected return, yet their primary use has been limited to risk-neutral settings. Conversely, existing risk-averse algorithms ensure safety by enforcing conservatism or restricted policy classes (Kumar et al., 2020; Urpí et al., 2021; Ma et al., 2021). This paper asks: Can we obtain safety without sacrificing expressiveness?
We answer in the affirmative by proposing the Risk-Aware Multimodal Actor-Critic (RAMAC) framework (Fig. 1). RAMAC couples an expressive generative actor with a distributional critic and differentiates a combination of behavioral cloning (BC) and distributional risk (instantiated with Conditional Value-at-Risk (CVaR)) gradients through the generative process (Di Castro et al., 2012; Chow et al., 2015), thereby unifying high expressiveness with robust tail-risk control, and reducing the out-of-distribution (OOD) action.
Prior offline-RL approaches can be organized by mechanism: (i) Policy regularization constrains the policy to the data manifold via divergence minimization or policy priors, improving stability but often sacrificing policy expressiveness on complex tasks with risk-neutral examples such as (Fujimoto et al., 2019; Wu et al., 2019; Kumar et al., 2019; Fujimoto & Gu, 2021) and risk-aware methods with prior-anchored perturbation designs such as (Urpí et al., 2021; Chen et al., 2024). (ii) Value conservatism reduces optimistic extrapolation, but can underestimate the value of infrequent yet high-return in-distribution modes due to global pessimism and data imbalance in both risk-neutral ((Kumar et al., 2020)) and risk-aware instances ((Ma et al., 2021)). (iii) Model-based pessimism bounds transition uncertainty with ensembles and penalties, at the cost of compounding model errors at scale again under both risk-neutral ((Yu et al., 2020; 2021; Rigter et al., 2022)) and risk-aware ((Rigter et al., 2023)) settings. (iv) Expressive generative policies faithfully clone multimodal behavior and achieve state-of-the-art mean returns, but limited use only in risk-neutral applications (Chen et al., 2021; Janner et al., 2022; Ajay et al., 2022; Wang et al., 2023; Hansen-Estruch et al., 2023; Park et al., 2025) including closest concurrent works pairing diffusion with distributional critics (Anonymous, 2025; Liu et al., 2025).
Despite compelling results from expressive models (e.g., diffusion, flow matching) in risk-neutral RL, their potential in offline risk-aware RL remains largely untapped.
Here, we aim to leverage the advantages of expressive policies without compromising risk-aversion or increasing the OOD action rate. To this end, inspired by the success of risk-neutral expressive policies such as  (Wang et al., 2023), RAMAC optimizes a joint objective composed of a BC and CVaR. The direct empirical BC term reduces BC estimation error effect observed in methods such as  (Nair et al., 2020) and hence, reduces the OOD visitation rate which is a critical issue in offline RL. The CVaR term in the objective removes risk-blindness (Fig. 2). We show that RAMAC yields high expected return while minimizing risk on complex multimodal offline benchmarks.
Our contributions can be summarized as:
  • Risk-aware expressive policy learning: We leverage expressive policies in the context of risk-aware RL and present two instantiations: RADAC (diffusion) and RAFMAC (flow matching).
  • Theoretical insight: We provide a theoretical discussion on how a BC regularized objective can improve performance in the context of offline RL. Driven by our theoretical results, we conjecture on one possible mechanism through which expressive policies can further improve performance.
  • Experimental evaluation: On the Stochastic-D4RL, our two instantiations (RADAC / RAFMAC) outperform baselines on CVaR while maintaining competitive mean return on most tasks.
Figure 1: RAMAC pipeline. From the offline buffer $\mathcal{D}$ (gray), the distributional critic $Z_{\phi}$ (green) fits the return law with a quantile loss and aggregates its lower tail into a CVaR signal. That signal is differentiated through the generative path of the actor $\pi_{\theta}$ (blue; diffusion or flow), which is trained with the composite objective $\mathcal{L}_{\pi}=\mathcal{L}_{\mathrm{BC}}+\eta\,\mathcal{L}_{\mathrm{Risk}}$ to shift mass away from low-quantile regions while staying on-manifold.
2 Preliminaries
Offline RL
We consider a finite-horizon Markov Decision Process (MDP) $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\gamma,H)$ (Sutton et al., 1998) and a fixed offline dataset $\mathcal{D}=\{(s_{t},a_{t},r_{t},s_{t+1})\}_{i=1}^{N}$ collected by an unknown behavior policy $\beta$ (Prudencio et al., 2023). Let $\operatorname{supp}(\mathcal{D})$ denote the dataset’s empirical state–action support. The objective is to learn a policy $\pi$ that maximizes the expected return $J(\pi)=\mathbb{E}_{\pi,P}[\sum_{t=0}^{H-1}\gamma^{t}r_{t}]$ without extra environment interaction. The central challenge is distributional shift (i.e., OOD): When $\pi$ visits $(s,a)\notin\operatorname{supp}(\mathcal{D})$, value estimates extrapolate and can be unsafe (Kumar et al., 2020). Prior work alleviates this issue with behavior regularization, conservative critics, or model-based pessimism.
Behavior–Regularized Actor–Critic (BRAC)
A large family of offline methods uses an actor–critic with an explicit proximity term to the behavior policy (Nair et al., 2020; Fujimoto & Gu, 2021; Wu et al., 2019; Kumar et al., 2019):
$\mathcal{L}_{\text{Actor}}(\theta)=\mathbb{E}_{s\sim\mathcal{D},\;a\sim\pi_{\theta}(\cdot\mid s)}\bigl[-\,Q_{\phi}(s,a)-\alpha\log\pi_{\theta}(a\mid s)\bigr],$ (1)
$\mathcal{L}_{\text{Critic}}(\phi)=\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D},\;a^{\prime}\sim\pi_{\theta}(\cdot\mid s^{\prime})}\bigl(Q_{\phi}(s,a)-[r+\gamma Q_{\bar{\phi}}(s^{\prime},a^{\prime})]\bigr)^{2}.$ (2)
We expand this variant into the behavior-regularized distributional actor-critic framework in which value term optimizes a coherent risk measure in place of the mean Q, yielding risk-aware updates while retaining the same BC regularize.
Distributional RL and Risk Measures
To move beyond expectation, distributional RL learns the return law via the distributional Bellman operator (Bellemare et al., 2017)
$(\mathcal{T}^{\pi}Z)(s,a)\;\stackrel{{\scriptstyle d}}{{=}}\;r(s,a)+\gamma\,Z(s^{\prime},a^{\prime}),\quad s^{\prime}\sim P(\cdot|s,a),\ a^{\prime}\sim\pi(\cdot|s^{\prime}).$ (3)
and often parameterizes the inverse CDF $Z_{\phi}(s,a;\tau)$ with an Implicit Quantile Network (IQN) (Dabney et al., 2018). Access to quantiles enables coherent risk measure $\mathcal{D}(\cdot)$, such as the CVaR. For a risk level $\alpha\in(0,1]$, the CVaR admits a dual form (Rockafellar et al., 2000):
$\mathrm{CVaR}_{\alpha}(X)\;=\;\inf_{\begin{subarray}{c}q\ll P,\,0\leq\frac{dq}{dP}\leq\frac{1}{\alpha}\end{subarray}}\mathbb{E}_{q}[X].$ (4)
and the integral form (used for actor gradients) is $\mathrm{CVaR}_{\alpha}(X)=\frac{1}{\alpha}\int_{0}^{\alpha}F_{X}^{-1}(\tau)\,d\tau$.
Expressive Generative Policies as Differentiable Trajectories
Recent actors generate an action by evolving a latent $z\!\sim\!\mathcal{N}(0,I)$ along a differentiable path (Janner et al., 2022; Wang et al., 2023). We focus on the two families:
(i) Diffusion policies follow a reverse-time SDE (Song et al., 2021b),
$\mathrm{d}_{t}\mathbf{a}_{t}=f_{\theta}(t,\mathbf{a}_{t},s)\,\mathrm{d}t+g(t)\,\mathrm{d}\mathbf{w}_{t},$ (5)
and (ii) Flow-matching policies solve a deterministic ODE (Lipman et al., 2023),
$\frac{\mathrm{d}\mathbf{a}_{t}}{\mathrm{d}t}=v_{\theta}(t,\mathbf{a}_{t},s).$ (6)
The entire map $\psi_{\theta}:s,z\mapsto a$ is differentiable, enabling guidance from a critic through denoising/integration steps. Prior work typically injects expected-value signals; our framework instead inject distributional risk signals.
3 Risk-Aware Multimodal Actor-Critic (RAMAC)
We now introduce the Risk-Aware Multimodal Actor-Critic (RAMAC). At its core, RAMAC operates in two stages: First, a distributional critic parameterized as an IQN learns the full conditional distribution of returns. Second, a generative actor, instantiated as either a diffusion policy or a flow-matching policy, is guided jointly by two terms in the objective function: (i) BC term that constrains the policy to the data manifold and (ii) CVaR term extracted from the critic’s lower tail. The latter pushes the probability mass away from low-probability, catastrophic regions and preserves the high-reward modes (Fig. 2). The former acts as a regularizer and limits OOD visitation rate.
3.1 Distributional Critic
Risk-sensitive objectives such as CVaR require access to the entire return distribution. We therefore adopt a distributional critic $Z_{\phi}$ via IQN (Dabney et al., 2018), building on the distributional Bellman operator in Eq. 3 Bellemare et al. (2017). We minimize a distributional Bellman residual with a quantile-Huber loss (with $\kappa\!=\!1$):
$\mathcal{L}_{\text{critic}}(\phi)=\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D},\;a^{\prime}\sim\pi_{\theta}(\cdot|s^{\prime}),\;\tau,\tau^{\prime}\sim\mathcal{U}(0,1)}\Big[\mathcal{L}_{\kappa}\!\big(r+\gamma Z_{\bar{\phi}}(s^{\prime},a^{\prime};\tau^{\prime})-Z_{\phi}(s,a;\tau)\,;\,\tau\big)\Big].$ (7)
This yields calibrated lower-tail quantiles that will directly drive the risk-aware actor update in Sec. 3.2.
Figure 2: RAMAC learning dynamics (conceptual). Top: policy density $\pi_{\theta}(a\!\mid\!s)$ induced by the reparameterized actor $a=\psi_{\theta}(s,z)$ (Eq. 8) over training. Bottom: critic return distribution $Z_{\phi}(s,a,\tau)$ with low quantiles highlighted (red); the actor is updated by the CVaR objective (Eqs. 911) while the critic is trained via the IQN loss (Eq. 7). CVaR updates steer mass away from low-quantile regions while preserving multimodal high-reward modes.
3.2 Risk-aware Generative Actor
An action is sampled as:
$a\;=\;\psi_{\theta}(s,z),\qquad z\sim\mathcal{N}(0,I).$ (8)
We define CVaR at level $\alpha$ through the critic’s quantiles and use a Monte Carlo estimator:
$\mathrm{CVaR}_{\alpha}\!\bigl(Z_{\phi}(s,a)\bigr)=\frac{1}{\alpha}\int_{0}^{\alpha}Z_{\phi}(s,a;\tau)\,d\tau\;\approx\;\frac{1}{K}\sum_{k=1}^{K}Z_{\phi}\bigl(s,a;\tau_{k}\bigr),\quad\tau_{k}\sim\mathcal{U}(0,\alpha).$ (9)
The risk loss maximizes this quantity. This is equivalent to minimizing the negative CVaR  [^1] This specific loss, instantiated with CVaR, is what we refer to as $\mathcal{L}_{\text{CVaR}}$ in our architectural diagrams for clarity.:
$\mathcal{L}_{\mathrm{Risk}}(\theta)=-\,\mathbb{E}_{s\sim\mathcal{D},\;a\sim\pi_{\theta}(\cdot|s)}\bigl[\mathrm{CVaR}_{\alpha}\!\big(Z_{\phi}(s,a)\big)\bigr].$ (10)
3.3 Behavior-regularized Objective
The complete policy objective balances risk aversion with fidelity to the offline dataset. It combines the risk term with a standard behavior cloning (BC) loss, $\mathcal{L}_{\text{BC}}(\theta)$:
$\mathcal{L}_{\pi}(\theta)=\underbrace{\mathcal{L}_{\text{BC}}(\theta)}_{\text{data fidelity}}+\eta\,\underbrace{\mathcal{L}_{\text{Risk}}(\theta)}_{\text{tail-risk aversion}}.$ (11)
where $\eta$ is a hyperparameter. In this work, we demonstrate both diffusion (RADAC) and flow (RAFMAC) variants. We show a pseudocode for RAMAC in Algorithm 1 and describe the full implementation details in App. C
Algorithm 1 Risk-Aware Multimodal Actor-Critic (RAMAC)
1:Initialize policy network $\pi_{\theta}$, critic $Z_{\phi}$, target critic $Z_{\bar{\phi}}$; mini–batch size $B$, risk level $\alpha$, critic–tail samples $K$, Exponential Moving Average (EMA) rate $\rho$.
2:repeat
3:  Sample a mini–batch $\{(s,a,r,s^{\prime})\}_{b=1}^{B}\sim\mathcal{D}$.
4:  Training Critic:
5:  Sample $z^{\prime}\!\sim\!\mathcal{N}(0,I)$ and set $a^{\prime}=\psi_{\theta}(s^{\prime},z^{\prime})$ (Eq. 8);
6:  Sample $\tau,\tau^{\prime}\!\sim\!\mathcal{U}(0,1)$ and update $\phi$ by minimizing $\mathcal{L}_{\text{critic}}(\phi)$ (Eq. 7).
7:  Training Actor:
8:  Sample $z\!\sim\!\mathcal{N}(0,I)$ and set $a=\psi_{\theta}(s,z)$ (Eq. 8);
9:  Sample $\tau_{1},\dots,\tau_{K}\!\sim\!\mathcal{U}(0,\alpha)$ and update $\theta$ by minimizing $\mathcal{L}_{\pi}(\theta)$ (Eq. 11).
10:  Target update: $\bar{\phi}\leftarrow\rho\,\bar{\phi}+(1-\rho)\,\phi$.
11:until converged
4 Behavior Regularization
Prior work has demonstrated the importance of behavior regularization in offline RL due to its ability to constrain the learned policy to the data manifold, curb value extrapolation, and stabilize improvement under distributional shift. A commonly adopted regularization scheme in offline risk-aware RL is prior-anchored perturbation method (e.g. ORAAC, UDAC) [^2] For simplicity and consistency with our experiments, we will refer to UDAC as ORAAC–Diffusion (Urpí et al., 2021; Chen et al., 2024), which uses a linear mixing of a pretrained BC policy with the RL actor (perturbation). Here, we first discuss the limitation of this regularization approach. We then demonstrate the advantages of our adopted scheme, namely, behavior-regularized objective method (shown in Eq. 11).
4.1 Prior-Anchored Perturbation and its limitations
In this approach, policy output can be written as:
$a\;=\;b\;+\;\zeta_{\psi}(s,b),\qquad b\sim G_{\phi}(\cdot\mid s),\ \ \|\zeta_{\psi}(s,b)\|\leq\Phi,$ (12)
where $\zeta_{\psi}$ is a learned residual (optimized to increase $Q$ or CVaR) and the norm bound $\Phi$ keeps updates closed to the anchor. Define the anchor support $\mathcal{S}_{G}(s)$ (the region in action space where $G_{\phi}(\cdot\mid s)$ places mass), the full action space $\mathbb{R}^{d}$, and the $\Phi$-radius ball of $b$
$B_{\Phi}(b)\;=\;\{\,a\in\mathbb{R}^{d}:\ \|a-b\|_{2}\leq\Phi\,\}.$
where any perturbed action $a=b+\zeta_{\psi}$ with $\|\zeta_{\psi}\|\leq\Phi$ lies in $B_{\Phi}(b)$. Hence on-manifold deployment is guaranteed by the safety margin condition
$\mathrm{dist}\!\big(b,\ \mathbb{R}^{d}\!\setminus\!\mathcal{S}_{G}(s)\big)\;>\;\Phi\quad\Longrightarrow\quad B_{\Phi}(b)\subseteq\mathcal{S}_{G}(s)\ \text{ and }\ a\in\mathcal{S}_{G}(s)\ \text{ for all }\|\zeta_{\psi}\|\leq\Phi,$
where $\mathrm{dist}(x,A)\!\mathrel{\mathop{:}}=\!\inf_{y\in A}\|x-y\|_{2}$ denotes Euclidean distance. OOD can still occur when this margin fails. This method provides a convenient local improvement rule, however, prior work has observed it suffers from poor mode coverage in multimodal action spaces (Wang et al., 2023). In addition to the identified limitations, we show distinct geometric weakness that can occur even without multimodality; having multiple modes merely magnifies the effect.
  • “Thin” support near $b$: “Thin” means the local supported region around $b$ is narrow; formally, the margin $m(b)\mathrel{\mathop{:}}=\mathrm{dist}\!\big(b,\mathbb{R}^{d}\!\setminus\!\mathcal{S}_{G}(s)\big)$ is small. If $m(b)\leq\Phi$, the ball $B_{\Phi}(b)$ overlaps the outside of $\mathcal{S}_{G}(s)$, so some $a=b+\zeta_{\psi}$ become OOD even though $\|\zeta_{\psi}\|\leq\Phi$.
  • Nonconvex support: “Nonconvex” means $\mathcal{S}_{G}(s)$ is not a convex set (e.g., a ring/annulus with a hole). Even if $b\in\mathcal{S}_{G}(s)$, a ball around $b$ may jump outside through a nearby concavity or hole whenever the margin $m(b)$ is small.
  • Gradient pushes off the data surface: The residual $\zeta_{\psi}(s,b)$ is trained to increase $Q$ or CVaR and is not constrained to be tangent to the data manifold. Consequently, gradients can point along the manifold’s normal direction, driving $a=b+\zeta_{\psi}$ toward the ball boundary ($\|\zeta_{\psi}\|\!\approx\!\Phi$). If $m(b)\!\leq\!\Phi$, these updates cross into OOD region.
4.2 Behavior‑Regularized Objective: Why it works better
In contrast to prior-anchored perturbation, as shown in Eq. 11, the BC term is applied directly to the deployed generative policy $\pi_{\theta}$. For explicit‑likelihood actors, the BC loss is the Negative Log-Likelihood (NLL): $\mathbb{E}_{(s,a)\sim\mathcal{D}}\!\left[-\log\pi_{\theta}(a\mid s)\right]=H(\beta(\cdot\mid s))+D_{\mathrm{KL}}\!\big(\beta(\cdot\mid s)\,\|\,\pi_{\theta}(\cdot\mid s)\big)$, so minimizing NLL shrinks the forward KL up to the data‑dependent constant $H(\beta)$. We therefore monitor the BC loss as a practical proxy for $D_{\mathrm{KL}}(\beta\|\pi_{\theta}).$
Proposition 1.
For each state $s$, let $I_{s}=\{a:\beta(a\mid s)>0\}$ and $O_{s}=I_{s}^{c}$. Assume $\beta\!\ll\!\pi_{\theta}$ (absolute continuity on $I_{s}$). Then the per-state OOD probability $\delta_{s}(\pi_{\theta})\mathrel{\mathop{:}}=\pi_{\theta}(O_{s}\mid s)$ satisfies
$\delta_{s}(\pi_{\theta})\ \leq\ 1-\exp\!\Big\{-D_{\mathrm{KL}}\big(\beta(\cdot\mid s)\,\|\,\pi_{\theta}(\cdot\mid s)\big)\Big\}.$ (13)
This shows that shrinking forward KL via BC can suppress per–state OOD, controlled by the selection of $\eta$ (Eq. 11), hence, avoiding an important challenge of offline RL discussed in Sec. 2 (Proof appears in App. A).
We can extend this observation to provide insight on one possible mechanism through which more expressive policies can support better performance in offline-RL settings. Consider the multi-modal data $D^{\text{M}}$, uni-modal policy, $\pi_{\omega}^{\text{U}}$, and multi-modal policy, $\pi_{\xi}^{\text{M}}$. We can conjecture that the following holds:
$\min_{\omega}\mathbb{E}_{(s,a)\sim\mathcal{D}^{\text{M}}}\!\left[-\log\pi^{\text{U}}_{\omega}(a\mid s)\right]\geq\ \min_{\xi}\mathbb{E}_{(s,a)\sim\mathcal{D}^{\text{M}}}\!\left[-\log\pi^{\text{M}}_{\xi}(a\mid s)\right]$
$\Rightarrow D_{\mathrm{KL}}(\beta\|\pi^{\text{U}}_{\omega^{*}})\geq\ D_{\mathrm{KL}}(\beta\|\pi^{\text{M}}_{\xi^{*}})$
As such, we can better effectuate lower OOD rate via Eq. 11 by adopting more expressive policies such as diffusion and flow-matching. It should be noted that the discussions here are agnostic to whether a risk-aware or risk-neutral formulation is adopted.
4.3 Example
We design a 2‑D contextual bandit with two disjoint modes (Toy Risky Bandit): Top left in Fig. 3 shows a ground truth that consists of safe center (moderate reward, no catastrophic tail) and a risky ring (higher mean with rare large penalties). The task isolates multimodality and lower-tail hazards. Below we introduce our baselines.
4.3.1 Expressive but risk‑neutral controls
(i) DiffusionQL (Wang et al., 2023): The actor is a conditional diffusion policy. An action is generated by the reverse path $a=\psi_{\theta}(s,z)$ (Eq. 5) with $z\!\sim\!\mathcal{N}(0,I)$. The BC term is the standard noise‑prediction (score‑matching) loss $L^{\text{BC}}_{\text{diff}}(\theta)=\mathbb{E}_{(s,a)\sim\mathcal{D},\ t,\ \varepsilon\sim\mathcal{N}}\!\bigl[\ \|\,\varepsilon-\varepsilon_{\theta}(s,\alpha_{t}a+\sigma_{t}\varepsilon,t)\|_{2}^{2}\ \bigr]$, and risk‑neutral improvement is applied by maximizing the critic’s expected value through the sampler, $L^{\text{Q}}_{\text{diff}}(\theta)\;=\;-\,\mathbb{E}_{s\sim\mathcal{D},\ z\sim\mathcal{N}}\bigl[\,Q_{\phi}\bigl(s,\psi_{\theta}(s,z)\bigr)\,\bigr]$. We use the combined objective $\mathcal{\min}_{\theta}\ L^{\text{BC}}_{\text{diff}}+\eta\,\mathcal{L}^{\text{Q}}_{\text{diff}}$.
(ii) FlowQL (Park et al., 2025): we maintain two policies. The BC flow policy $\mu_{\theta}(s,z)$ is trained only with the flow-matching BC loss (Eq. 6). Alongside, we train a one-step policy $\mu_{\omega}(s,z)$ with the actor loss $L_{\pi}^{\text{FQL}}(\omega)=\mathcal{L}^{\text{distill}}_{\text{flow}}+\eta\,\mathcal{L}^{\text{Q}}_{\text{flow}}$, and deploy $\mu_{\omega}$ at test time (no iterative flow needed).
(iii) Conditional VAE(CVAE)-QL: An autoregressive conditional decoder parameterizes $p_{\theta}(a\mid s)=\prod_{i=1}^{d}p_{\theta}(a_{i}\mid a_{<i},s)$. BC is the NLL $\mathcal{L}^{\text{BC}}_{\text{cvae}}=-\mathbb{E}_{\mathcal{D}}\log p_{\theta}(a\mid s)$, and the risk‑neutral improvement is the same.
4.3.2 Prior‑anchored perturbation (risk‑aware)
This category of baselines adopt Eq. 12:
(i) ORAAC (Urpí et al., 2021): samples an anchor $b\!\sim\!G_{\phi}(\cdot\mid s)$ from a behavior prior and applies a bounded perturbation toward the actor while optimizing a coherent risk objective (e.g., CVaR);
(ii) ORAAC–Diffusion (Chen et al., 2024): replaces the VAE prior with a diffusion prior, keeping the same anchor–perturb structure;
(iii)ORAAC–Flow: a flow‑prior counterpart in the same mixing form. These (B) baselines let us test the geometric leakage.
4.3.3 Example Results
The resulting policy distributions for various methods are shown in Fig. 3.
Risk‑neutral expressive controls (Fig. 3 b-d): Overall, as expected, these methods are risk-blind and they chase high‑$Q$ ridges without regard to the lower tail. FlowQL often preserves both modes but does not suppress mass on the hazardous ring; Diffusion QL drifts toward sparsely covered high‑$Q$ pockets on the ring, yielding risk exposure; the CVAE variant collapses topology and fills low‑density bridges.
Prior‑anchored perturbation (Fig. 3 e-g): ORAAC and its diffusion/flow variants place substantial mass in the inter‑mode low‑density region, regardless of whether bc prior is expressive or not.
RADAC (Fig. 3 h): by sending CVaR signals from a distributional critic through the diffusion/flow trajectory while regularizing with BC, RADAC concentrates probability near the safe center without filling the gap. Full configuration and additional plots are in App. E.1.
(a) Ground Truth
(b) CVAE-QL
(c) DiffusionQL
(d) FlowQL
(e) ORAAC
(f) ORAAC-Diffusion
(g) ORAAC-Flow
(h) RADAC
Figure 3: Toy Risky Bandit Results Top: Ground Truth consists of a safe center mode yellow-green and a risky ring where high-reward samples yellow are interspersed with catastrophic penalties (purple). Risk-neutral generative baselines concentrate on the risky ring or collapses topology. Bottom: Prior‑anchored perturbation methods produce samples in the low‑density inter‑mode region, exhibiting OOD leakage. RADAC concentrates near the safe center without losing multimodality. See App. D for more results.
5 Experiments
In this section, we evaluate RADAC and RAFMAC on the Stochastic-D4RL benchmarks to validate both risk awareness and policy expressiveness. We also quantify the OOD action rate $\varepsilon_{\mathrm{act}}$ (Sec. 5.3) to link practice to our view (Sec 4). Additional results appear in App. D.
Tasks
We augment standard D4RL locomotion tasks (Fu et al., 2020) with rare heavy‑tailed penalties tied to velocity or torso pitch angles with early termination following  (Urpí et al., 2021) (full construction details and per‑task parameters in App. E). We evaluate on hopper, walker2d, and halfcheetah using the medium–expert and medium–replay datasets, which are multimodal by construction (mixtures of heterogeneous behaviors). It lets us examine that RAMAC learns risk‑aware policies without sacrificing multimodality. During training, we follow  (Urpí et al., 2021) and relabel per-transition rewards in the offline datasets with the same stochastic hazard model; evaluation uses the identical hazard specification.
Baselines
We compare against representative offline‑RL methods covering value conservatism, distributional conservatism, anchor–perturb risk aversion, and risk–neutral expressive generators: CQL (Kumar et al., 2020), CODAC (Ma et al., 2021), ORAAC (Urpí et al., 2021), DiffusionQL (Wang et al., 2023), and FlowQL (Park et al., 2025).
Evaluation
Following protocols as those adopted in  (Urpí et al., 2021; Wang et al., 2023), we train for 2000 epochs, each with 1000 gradient steps and batch size 256. We evaluate methods at fixed intervals of gradient steps and report (i) raw returns averaged over 5 seeds and (ii) episodic $\mathrm{CVaR}_{0.1}$ computed over 50 rollouts in total (10 evaluation episodes per seed) to avoid normalization bias on the stochastic variants. For ORAAC and CODAC, we adopt the authors’ risk-aware objectives (risk level $\alpha{=}0.1$ unless noted). For the other baselines, we tune hyperparameters within the same training budget to ensure fairness and otherwise use authors’ recommended settings (App. E.3). Further protocol details appear in App. E. For artifact reproducibility, following common practice, we provide full results with corresponding 1000-step evaluation in App. D.
5.1 Results and Analysis
Table 1 reports Mean and CVaR0.1 for RADAC and RAFMAC alongside baselines. Across six tasks, both RAMAC instantiations deliver strong lower tails with competitive or higher means. Viewed through mode coverage, the diffusion actor in RADAC updates actions over small denoising steps, which tends to maintain diverse in–support modes while letting CVaR guidance move probability away from hazardous boundaries (Dhariwal & Nichol, 2021); this explains RADAC’s stronger CVaR performance on halfcheetah and hopper‑medium‑replay under knife-edge hazards. The flow-matching actor in RAFMAC transports along a short, deterministic ODE path that efficiently sharpens dominant high-reward modes (Lipman et al., 2023) on walker2d‑medium‑expert, often lifting mean return when hazards are smoother. ORAAC regularizes toward a behavior anchor. It reliably handles sharp hazardous thresholds on such as hopper‑medium‑expert but may fail to exploit high-reward modes, and can place mass in low-density between-mode regions depending on the anchor–perturbation. These tendencies align with qualitative safety plots (Fig. 4).
Table 1: Stochastic–D4RL results over 5 seeds. We report Mean and CVaR0.1; best in dark/ second in light shaded. The Full results with s.e. appear in App. D.2.
Dataset Metric CQL CODAC ORAAC FlowQL DiffusionQL RAFMAC RADAC
HalfCheetah-medium-expert Mean 66.66-66.66 0.12-0.12 $796.06$ $844.14$ 20.71-20.71 $889.56$ $\mathbf{916.64}$
CVaR 135.39-135.39 0.11-0.11 $742.94$ $754.44$ 76.39-76.39 $736.95$ $\mathbf{805.25}$
Walker2d-medium-expert Mean 21.52-21.52 $23.96$ $969.62$ $1309.48$ 32.38-32.38 $\mathbf{1822.24}$ $1708.68$
CVaR 64.88-64.88 43.88-43.88 $358.55$ $468.15$ 116.19-116.19 $\mathbf{1127.21}$ $573.22$
Hopper-medium-expert Mean 25.87-25.87 $26.59$ $\mathbf{714.15}$ $341.16$ 279.97-279.97 $281.24$ $130.74$
CVaR 111.37-111.37 150.92-150.92 $\mathbf{374.63}$ 8.80-8.80 872.95-872.95 132.33-132.33 167.29-167.29
HalfCheetah-medium-replay Mean 66.21-66.21 0.11-0.11 $18.99$ $434.33$ $279.95$ $449.04$ $\mathbf{525.84}$
CVaR 127.09-127.09 1.47-1.47 34.09-34.09 $224.73$ $79.93$ $144.73$ $\mathbf{278.65}$
Walker2d-medium-replay Mean 16.90-16.90 $33.59$ $126.94$ $411.36$ $96.88$ 71.69-71.69 $\mathbf{615.94}$
CVaR 51.49-51.49 52.63-52.63 203.64-203.64 $5.08$ $48.14$ $\mathbf{530.37}$ $145.21$
Hopper-medium-replay Mean 16.25-16.25 47.83-47.83 18.00-18.00 $373.16$ 2.79-2.79 $303.44$ $\mathbf{385.58}$
CVaR 118.70-118.70 160.08-160.08 129.25-129.25 62.24-62.24 51.33-51.33 90.73-90.73 $\mathbf{-8.16}$
5.2 Qualitative Safety Analysis
We visualize a three representative method’s contrast, risk-aware expressive generator RADAC, risk-neutral expressive generator DiffusionQL, and anchor-perturb risk-averse method ORAAC. Fig. 4 plots the monitored distribution of policies against safe regions. RADAC concentrate probability mass inside or near safe boundary while actively reallocates probability onto high‑return modes that lie within the safety regions. DiffusionQL is tightly concentrated around zero because rare, high penalties depress bootstrapped values near the safe boundary. On the other hand, ORAAC regularizes toward a behavior anchor and thus settles at a low density between‑mode when the anchor lies in risky-region. These features can be further observed in Table 8 in App. D.2
(a) Halfcheetah-Medium-Replay
Density Refer to caption05
(b) Walker2d-Medium-Replay
Refer to caption-0.500.5
(c) Hopper-Medium-Replay
Refer to caption-0.100.1
(d) Halfcheetah-Medium-Expert
Density Refer to caption0510
(e) Walker2d-Medium-Expert
Refer to caption-0.500.5
(f) Hopper-Medium-Expert
Refer to caption-0.100.1
Figure 4: Policy distributions for RADAC, ORAAC, and DiffusionQL; shaded bands indicate safe operational ranges (HalfCheetah: $v\!\leq\!10$ for m-e, $v\!\leq\!5$ for m-r; Hopper: $|\theta|\!\leq\!0.1$; Walker2d: $|\theta|\!\leq\!0.5$). RADAC reduces mass beyond thresholds.
5.3 Empirical Analysis of Theoretical Insights
We now provide measurements of OOD actions to validate the insights in Sec. 4. For each policy, we report $\varepsilon_{\text{act}}$, the fraction of evaluation actions whose 1-NN distance to the dataset exceeds $\kappa\!\times\!\mathrm{median}\,d_{\mathrm{NN}}$. Sec. 5.2 predicts that (i) (expressive, BC-regularized CVaR objective) should reweight probability within the data manifold, yielding low $\varepsilon_{\text{act}}$. (ii) ORAAC, being less expressive and based on anchor–perturbation, should exhibit higher $\varepsilon_{\text{act}}$ than RADAC; Table 2 confirms this prediction: RADAC retains low-OOD across task, consistent with BC-regularized OOD suppression; ORAAC is consistently higher than RADAC, matching the expected geometric leakage from anchor–perturbation, consistent with the Sec. 4. RADAC achieves risk awareness and expressiveness simultaneously with low-OOD.
Table 2: OOD action rate ($\%\pm$ s.e.) on medium–expert. $\kappa{=}3$).
Task RADAC (ours) ORAAC
HalfCheetah $2.04\!\pm\!0.80$ $6.15\!\pm\!1.5$
Walker2d $0.75\!\pm\!0.54$ $10.84\!\pm\!1.98$
Hopper $0.77\!\pm\!0.56$ $2.68\!\pm\!1.01$
6 Conclusions
This paper introduces RAMAC, a model-free framework for risk-aware offline RL using expressive generative policies. This is done by coupling a distributional critic with diffusion/flow actors and simultaneously incorporating a CVaR together with a BC regularization component in the objective function. We provide a theoretical explanation and proof on how BC regularization through the objective can reduce OOD action rate. This observation further provides insight on one possible mechanism through which expressive policies such as diffusion and flow can be effective in the context of offline RL. We confirm our observation through an example and finally, evaluate the method against several baselines on the Stochastic-D4RL benchmark. Our proposed approach improves CVaR on most tasks while maintaining competitive mean return.
Appendix A Proof of Proposition 1
Proof.
Fix a state $s$ and define the $\beta$–support $I_{s}\mathrel{\mathop{:}}=\{a:\beta(a\mid s)>0\}$ and its complement $O_{s}\mathrel{\mathop{:}}=I_{s}^{c}$. Assume $\beta\ll\pi_{\theta}$ on $I_{s}$ (so $\pi_{\theta}(I_{s}\mid s)>0$ and the forward KL is finite).
Since $\beta(\cdot\mid s)$ has all its mass on $I_{s}$,
$D_{\mathrm{KL}}(\beta\|\pi_{\theta})$ $=\int_{\mathcal{A}}\beta(a\mid s)\,\log\frac{\beta(a\mid s)}{\pi_{\theta}(a\mid s)}\,da\;=\;\int_{I_{s}}\beta(a\mid s)\,\log\frac{\beta(a\mid s)}{\pi_{\theta}(a\mid s)}\,da.$ (1)
Define the normalization of $\pi_{\theta}$ to $I_{s}$:
$\pi_{I}(a\mid s)\;\mathrel{\mathop{:}}=\;\pi_{\theta}(a\mid s,\ a\in I_{s})\;=\;\frac{\pi_{\theta}(a\mid s)}{\pi_{\theta}(I_{s}\mid s)},\qquad a\in I_{s},$
so that on $I_{s}$ we have the identity $\pi_{\theta}(a\mid s)=\pi_{\theta}(I_{s}\mid s)\,\pi_{I}(a\mid s)$.
Substitute the above factorization into (1) and use $\log(xy)=\log x+\log y$:
$\log\frac{\beta(a\mid s)}{\pi_{\theta}(a\mid s)}$ $=\log\frac{\beta(a\mid s)}{\pi_{\theta}(I_{s}\mid s)\,\pi_{I}(a\mid s)}=\log\frac{\beta(a\mid s)}{\pi_{I}(a\mid s)}-\log\pi_{\theta}(I_{s}\mid s).$ (2)
Plug (2) into (1) and split the integral:
$D_{\mathrm{KL}}(\beta\|\pi_{\theta})$ $=\int_{I_{s}}\beta(a\mid s)\,\log\frac{\beta(a\mid s)}{\pi_{I}(a\mid s)}\,da\;-\;\int_{I_{s}}\beta(a\mid s)\,\log\pi_{\theta}(I_{s}\mid s)\,da$
$=\underbrace{D_{\mathrm{KL}}\big(\beta(\cdot\mid s)\,\|\,\pi_{I}(\cdot\mid s)\big)}_{\geq 0}\;-\;\log\pi_{\theta}(I_{s}\mid s)\,\underbrace{\int_{I_{s}}\beta(a\mid s)\,da}_{=1}.$ (3)
Here the last equality uses that $\log\pi_{\theta}(I_{s}\mid s)$ is constant in $a$, and that $\beta$ places total mass $1$ on $I_{s}$.
From (3) and nonnegativity of KL,
$D_{\mathrm{KL}}(\beta\|\pi_{\theta})\;\geq\;-\,\log\pi_{\theta}(I_{s}\mid s).$
Exponentiating both sides gives
$e^{-D_{\mathrm{KL}}(\beta\|\pi_{\theta})}\;\leq\;\pi_{\theta}(I_{s}\mid s).$
Since $\pi_{\theta}(I_{s}\mid s)=1-\delta_{s}(\pi_{\theta})$ with $\delta_{s}(\pi_{\theta})\mathrel{\mathop{:}}=\pi_{\theta}(O_{s}\mid s)$, we obtain the per–state OOD bound
$\delta_{s}(\pi_{\theta})\ \leq\ 1-\exp\!\bigl\{-D_{\mathrm{KL}}(\beta(\cdot\mid s)\|\pi_{\theta}(\cdot\mid s))\bigr\}.\qed$
Appendix B Related Works
We review works most relevant to our risk-aware generative trajectory view—policies that map noise to actions through a differentiable path and how safety is enforced therein—while avoiding repetition of the core background already covered in the main text. For a broad taxonomy of offline RL, see Prudencio et al. (2023).
Expressive generative policies
The main paper reviews diffusion and flow-matching policies (e.g., DiffusionQL, Diffuser, IDQL, FlowQL). Here we note complementary developments not detailed there: (i) DDIM-style imitation learning that accelerates inference while keeping diffusion’s expressiveness (Song et al., 2021a); (ii) real-robot deployments of diffusion policies demonstrating hardware viability (Chi et al., 2023); and on the flow side . These works bolster the case for expressive, differentiable policies but remain risk-neutral in objective design.
Autoregressive generative baselines
Trajectory Transformer (Janner et al., 2021) provides strong risk-neutral baselines by modeling returns/actions autoregressively; diffusion has also been used for open-loop planning (Chen et al., 2023). Because decoding is single-shot, these approaches lack a continuous generative path through which tail-risk gradients can be injected, leading to high mean performance without explicit lower-tail control.
Risk-sensitive RL
Beyond expectation-oriented objectives, risk-sensitive control formalizes tail-aware criteria via coherent/dynamic risk measures for MDPs (Ruszczyński, 2010). Among coherent measures, CVaR admits sampling- and policy-gradient formulations suitable for RL (Tamar et al., 2015b; a), and has been linked to robustness via CVaR–robust trade-offs (Chow et al., 2015). In the offline regime, safety is often operationalized as high-confidence off-policy evaluation/improvement from fixed logs—e.g., HCOPE/HCPI and SPIBB (Thomas et al., 2015; Laroche et al., 2019)—which bound deployment risk yet do not address how expressive generators should receive lower-tail gradients.
Closest lines and delineation
Concurrent actor–critic lines that couple diffusion with value learning remain expectation-oriented: (Anonymous, 2025) stabilizes online diffusion actors with distributional critics and double-$Q$ but does not backpropagate CVaR along the denoising path ;(Fang et al., 2025)formulates offline constrained policy iteration as diffusion noise regression under KL/BC regularization ; and (Ma et al., 2025) studies efficient online diffusion control from an energy-based perspective. Distributional SAC variant (Ma et al., 2020) improve risk sensitivity via value-law estimation—typically with Gaussian policies—yet still lack CVaR shaping ; the diffusion-policy instantiation (Liu et al., 2025) targets multi-modality but likewise reports no CVaR along the multi-step generation . Risk-averse offline methods relying on behavior priors—e.g.(Urpí et al., 2021) ,and diffusion-prior (Chen et al., 2024)—use anchor–perturb/mixing mechanisms , while (Ma et al., 2021) imposes conservative distributional critics (value pessimism) . These approaches either (i) optimize expectation-oriented objectives with expressive generators or (ii) control risk via mixing/pessimism, in contrast to our distributional risk shaping without anchor mixing.
Appendix C Implementation Details
Actor architecture
RAMAC employs a reparameterized generative actor $a=\psi_{\theta}(s,z)$ so that gradients from the risk term flow through the entire generative trajectory. RADAC instantiates $\psi_{\theta}$ as a denoising diffusion policy with VP schedule and T=5T{=}5 denoising steps; the score network is an MLP (hidden 256–256, SiLU). RAFMAC instantiates $\psi_{\theta}$ as a deterministic flow–matching ODE solved by Euler with flow_steps K=10K{=}10; the velocity field is an MLP (hidden 256–256, SiLU). For both, the actor objective is $\mathcal{L}_{\pi}=\lambda_{\mathrm{BC}}\mathcal{L}_{\mathrm{BC}}+\eta\,\mathcal{L}_{\mathrm{Risk}}$, where $\mathcal{L}_{\mathrm{BC}}$ is the model’s native BC loss (score matching for diffusion, velocity matching for flow), and $\mathcal{L}_{\mathrm{Risk}}=-\mathbb{E}_{s,a\sim\pi_{\theta}}[\mathrm{CVaR}_{\alpha}(Z_{\phi}(s,a))]$ with $\alpha{=}0.1$.
Distributional critic architecture
Both variants share a Double IQN critic trained with the quantile Huber loss ($\kappa{=}1$). Two critics $Z_{\phi_{1}},Z_{\phi_{2}}$ are updated against a min target to curb overestimation; quantiles for TD use $\tau,\tau^{\prime}\!\sim\!\mathcal{U}(0,1)$, while the actor’s CVaR term samples $\tau\!\sim\!U(0,\alpha)$.
For a batch $(s,a,r,s^{\prime})\!\sim\!\mathcal{D}$, we generate the bootstrapping action via the actor:
$a^{\prime}\;=\;\psi_{\theta}(s^{\prime},z^{\prime}),\qquad z^{\prime}\sim\mathcal{N}(0,I),$
so that gradients (later used for risk shaping) can flow through the full generative trajectory (short reverse diffusion for RADAC; short ODE flow for RAFMAC). Right after specifying RADAC/RAFMAC actor parameters, we clarify the pre-loss stage for the critic before introducing the final loss. Instead of sampling $\tau$, we use a fixed uniform grid
$\mathcal{T}_{N}\;=\;\Big\{\,\tau_{i}\;=\;\tfrac{i-\tfrac{1}{2}}{N}\,\Big\}_{i=1}^{N}.$ (14)
For CVaR at level $\alpha$, let $m\!=\!\lfloor\alpha N\rfloor$; then
$\mathrm{CVaR}_{\alpha}\!\big(Z_{\phi}(s,a)\big)\;\approx\;\frac{1}{m}\sum_{i=1}^{m}Z_{\phi}\!\big(s,a;\tau_{i}\big),\qquad\tau_{i}\in\mathcal{T}_{N}.$ (15)
This is equivalent in expectation to drawing $\tau\!\sim\!\mathcal{U}(0,\alpha)$ (cf. Eq. 9) but with lower estimator variance.
We form target quantiles on another grid $\mathcal{T}_{N^{\prime}}$ and define the TD residual
$\delta_{\tau_{i},\tau^{\prime}_{j}}\;=\;r+\gamma\,Z_{\bar{\phi}}(s^{\prime},a^{\prime};\tau^{\prime}_{j})\;-\;Z_{\phi}(s,a;\tau_{i}),\qquad\tau_{i}\!\in\!\mathcal{T}_{N},\ \tau^{\prime}_{j}\!\in\!\mathcal{T}_{N^{\prime}}.$
With this pre-loss construction, the final critic objective is exactly the quantile-Huber residual minimization in Eq. 7/17; determinism only replaces stochastic $(\tau,\tau^{\prime})$ by $(\tau_{i},\tau^{\prime}_{j})$ from fixed grids.
The critic minimises the quantile-Huber loss (Dabney et al., 2018; Rowland et al., 2019)
$\mathcal{L}_{\kappa}(\delta;\tau)=\bigl|\tau-\mathbf{1}_{\{\delta<0\}}\bigr|\times\begin{cases}\tfrac{\delta^{2}}{2\kappa},&|\delta|\leq\kappa,\\[4.0pt] |\delta|-\tfrac{\kappa}{2},&\text{otherwise},\end{cases}$ (16)
with $\kappa\!=\!1$. Averaging over $N\times N^{\prime}$ quantile pairs yields
$\mathcal{L}_{\text{critic}}(\phi)=\mathbb{E}_{(s,a,r,s^{\prime}),\,a^{\prime}}\!\biggl[\frac{1}{NN^{\prime}}\!\sum_{i=1}^{N}\sum_{j=1}^{N^{\prime}}\mathcal{L}_{\kappa}\bigl(\delta_{\tau_{i},\tau^{\prime}_{j}};\,\tau_{i}\bigr)\biggr].$ (17)
Optimising Eq. 17 yields a calibrated estimate of the return law, whose lower tail supplies the CVaR gradients used in Step 2 (Sec. 3.2).
Hyperparameters
Unless noted, we use Adam for all networks (default $3{\times}10^{-4}$), batch size 256, discount $\gamma{=}0.99$, soft target update $\tau_{\text{target}}{=}0.005$, and no LR decay. RAMAC’s (critic LR, IQN size, $\eta$, gradient–norm clipping, optional QQ–target clipping, etc.) are listed in Table 6.
RAFMAC risk weight tuning
we swept $\eta\in\{1,10,50,100,300,1000\}$ and unified to $\eta{=}1000$ for all datasets; critic settings are fixed ($lr_{\text{critic}}{=}3{\times}10^{-4}$, emb_dim=128=128, n_quantiles=32=32) (Table 6).
Critic–target clipping
Where specified , target returns are clipped ([300,300][-300,300] or [150,150][-150,150]) to dampen rare outliers without affecting on–manifold learning.
Appendix D Additional Experimental Results
D.1 More 2D Synthetic Task Results
Behavior cloning task Fig. 5
On the 2D bandit dataset, three BC models show generator-specific patterns. CVAE-BC collapses topology and places probability in the low‑density gap. Diffusion-BC most faithfully reproduces both the ring and the inner cluster with appropriate thickness. Flow-Matching BC renders draws a sharp ring but allocates less mass to the center and shows edges spread slightly outward.
RADAC dynamics over training Fig. 6
Epoch-by-epoch samples show how the CVaR term reshapes a diffusion policy under BC. Early iterations spread mass over both modes; by $\sim$200 epochs the policy starts vacating the ring; Between 400 and 800 epochs the ring thins and probability shifts inward, while the central cluster grows; By roughly 950 epochs most mass is at the safe center. The final plot at epoch=1000 is in Fig. 3 Throughout, BC keeps samples on-manifold, so lower-tail risk is reduced without collapsing mode.
More toy results Fig. 7
shows the qualitative pattern is consistent across seeds. CVAE‑QL fills the low‑density gap; Diffusion‑QL and Flow-QL stay on the ring (high mean, higher risk), and anchor–perturb variants (ORAAC family) place samples in the inter-mode region. RADAC reassigns almost all probability to the safe center, whereas RAFMAC leaves a thinner ring. We attribute this to geometry: flow matching transports density via a smooth, near‑invertible ODE field, which preserves shape and favors thinning rather than removing the ring under BC; diffusion uses a stochastic reverse process whose stepwise CVaR guidance can reallocate mass across the low‑density gap.
(a) Ground Truth
(b) CVAE-BC
(c) Diffusion-BC
(d) Flow Matching-BC
Figure 5: Behavior cloning on the Risk Bandit dataset. Each panel shows i.i.d. samples from the BC Policy. CVAE‑BC mixes modes and places points in the low‑density gap; Diffusion‑BC reproduces both the outer ring and the central cluster; Flow‑Matching BC yields a crisp ring but assigns less mass to the center.
(a) epoch=0
(b) epoch=200
(c) epoch =400
(d) epoch=600
(e) epoch=800
(f) epoch=950
Figure 6: RADAC dynamics on the toy task. Mass gradually moves from the risky ring to the safe center: the ring thins (400–800 epochs) and the central cluster grows, ending with most mass at the center ( 950 epochs). BC keeps the policy on-manifold while CVaR reduces lower-tail risk.
Figure 7: More toy results. The qualitative pattern is unchanged across seeds; we show another seed. RAFMAC’s thinner ring and inward shift arise from CVaR‑shaped flow transport under BC.
[…]
Environment, Dataset Algorithm Mean CVaR
HalfCheetah-m-e CQL $66.66\!\pm\!13.17$ $135.39\!\pm\!27.71$
CODAC $2\!\pm\!0.16 0.11±0$$1\!\pm\!0.25$
ORAAC $796.06\!\pm\!30.28$ $742.94\!\pm\!22.95$
FlowQL $844.14\!\pm\!16.15$ $754.44\!\pm\!27.26$
DiffusionQL $20.71\!\pm\!18.89$ $76.39\!\pm\!14.39$
RAFMAC $889.56\!\pm\!38.31$ $736.95\!\pm\!102.54$
RADAC $\mathbf{916.64}\!\pm\!35.80$ $\mathbf{805.25}\!\pm\!15.34$
Walker2d-m-e CQL $21.52\!\pm\!8.68$ $64.88\!\pm\!18.32$
CODAC $23.96\!\pm\!10.56$ $43.88\!\pm\!13.28$
ORAAC $969.62\!\pm\!442.36$ $358.55\!\pm\!682.29$
FlowQL $1309.48\!\pm\!233.72$ $468.15\!\pm\!416.61$
DiffusionQL $32.38\!\pm\!64.15$ $116.19\!\pm\!46.39$
RAFMAC $\mathbf{1822.24}\!\pm\!128.36$ $\mathbf{1127.21}\!\pm\!620.95$
RADAC $1708.68\!\pm\!163.19$ $573.22\!\pm\!894.62$
Hopper-m-e CQL $25.87\!\pm\!13.46$ $111.37\!\pm\!51.69$
CODAC $26.59\!\pm\!47.56$ $150.92\!\pm\!42.19$
ORAAC $\mathbf{714.15}\!\pm\!243.57$ $\mathbf{374.63}\!\pm\!326.66$
FlowQL $341.16\!\pm\!75.98$ $8.80\!\pm\!84.57$
DiffusionQL $279.97\!\pm\!215.46$ $872.95\!\pm\!589.90$
RAFMAC $281.24\!\pm\!82.07$ $132.33\!\pm\!183.92$
RADAC $130.74\!\pm\!273.53$ $167.29\!\pm\!107.33$
HalfCheetah-m-r CQL $66.21\!\pm\!11.52$ $127.09\!\pm\!37.10$
CODAC $0.11\!\pm\!0.16$ $1.47\!\pm\!0.53$
ORAAC $18.99\!\pm\!34.67$ $34.09\!\pm\!25.47$
FlowQL $434.33\!\pm\!40.45$ $224.73\!\pm\!146.83$
DiffusionQL $279.95\!\pm\!91.48$ $79.93\!\pm\!110.85$
RAFMAC $449.04\!\pm\!73.84$ $144.73\!\pm\!181.54$
RADAC $\mathbf{525.84}\!\pm\!44.61$ $\mathbf{278.65}\!\pm\!151.27$
Walker2d-m-r CQL $16.90\!\pm\!7.56$ $51.49\!\pm\!14.17$
CODAC $33.59\!\pm\!45.29$ $52.63\!\pm\!42.63$
ORAAC $126.94\!\pm\!178.91$ $203.64\!\pm\!338.87$
FlowQL $411.36\!\pm\!70.84$ $5.08\!\pm\!240.85$
DiffusionQL $96.88\!\pm\!198.31$ $48.14\!\pm\!227.71$
RAFMAC $71.69\!\pm\!241.69$ $\mathbf{530.37}\!\pm\!84.57$
RADAC $\mathbf{615.94}\!\pm\!219.44$ $145.21\!\pm\!39.43$
Hopper-m-r CQL $6.25\!\pm\!20.60 11$$18.70\!\pm\!106.89$
CODAC $47.83\!\pm\!32.01$ $160.08\!\pm\!60.90$
ORAAC $18.00\!\pm\!44.92$ $129.25\!\pm\!108.63$
FlowQL $373.16\!\pm\!109.86$ $62.24\!\pm\!203.02$
DiffusionQL $2.79\!\pm\!12.83$ $51.33\!\pm\!36.90$
RAFMAC $303.44\!\pm\!28.95$ $90.73\!\pm\!93.82$
RADAC $\mathbf{385.58}\!\pm\!55.20$ $\mathbf{-8.16}\!\pm\!92.79$
D.2 Extended Stochastic-D4RL Results
Protocol
To remove post-hoc checkpoint selection and ease reproducibility, we report a full result in Sec. 3.1 with s.e. in Table 3 and fixed 1000-gradient-step evaluation for every method and task in Table 4. Scores are raw returns and episodic $\mathrm{CVaR}_{0.1}$ (mean $\pm$ s.e. over 5 seeds), without normalization, matching the stochastic variants used in the main text.
Consistency with the main-text trends
At the fixed 1000–step evaluation, the ranking patterns largely match the main text, but the mechanisms are task–dependent. Flow–based policies (FlowQL/RAFMAC) often reach higher mean by 1000 steps because flow matching uses a deterministic ODE with a short generative path and low-variance policy gradients; combined with velocity-matching BC, this yields fast on-manifold improvement. CVaR outcomes depend on lower-tail calibration of the distributional critic: with smooth, non-terminating penalties (e.g., HalfCheetah) RADAC/RAFMAC already improve CVaR at 1000 steps, whereas with sparse, terminating hazards (e.g., Hopper) ORAAC’s anchor regularization provides more stable early CVaR and mean. Walker2d sits in between: RAFMAC attains the highest mean at 1000 steps, and CVaR leadership alternates between FlowQL and RAFMAC depending on the dataset variant.
[…]
Environment, Dataset Algorithm Mean CVaR
HalfCheetah-m-e CQL $0.97\!\pm\!0.24$ $2.24\!\pm\!0.43$
CODAC $0.12\!\pm\!0.08$ $1.48\!\pm\!0.27$
ORAAC $4106.25\!\pm\!177.48$ $3692.79\!\pm\!466.31$
FlowQL $4695.46\!\pm\!65.97$ $4025.12\!\pm\!230.08$
DiffusionQL $118.72\!\pm\!64.53$ $198.01\!\pm\!76.76$
RAFMAC $5084.12\!\pm\!230.43$ $3735.37\!\pm\!827.60$
RADAC $5659.40\!\pm\!131.94$ $4667.96\!\pm\!42.59$
Walker2d-m-e CQL $10.32\!\pm\!6.27$ $73.38\!\pm\!9.02$
CODAC $27.56\!\pm\!6.26$ $35.30\!\pm\!15.36$
ORAAC $663.23\!\pm\!181.31$ $205.21\!\pm\!65.45$
FlowQL $2457.68\!\pm\!208.80$ $448.48\!\pm\!208.81$
DiffusionQL $32.33\!\pm\!4.59$ $68.43\!\pm\!11.28$
RAFMAC $3567.89\!\pm\!206.63$ $356.20\!\pm\!987.34$
RADAC $2760.21\!\pm\!689.32$ $322.76\!\pm\!757.44$
Hopper-m-e CQL $43.22\!\pm\!29.48$ $65.90\!\pm\!36.42$
CODAC $31.59\!\pm\!28.74$ $77.88\!\pm\!34.33$
ORAAC $660.07\!\pm\!157.55$ $400.84\!\pm\!142.60$
FlowQL $393.64\!\pm\!27.75$ $77.93\!\pm\!60.53$
DiffusionQL $38.75\!\pm\!27.68$ $212.49\!\pm\!91.99$
RAFMAC $370.11\!\pm\!39.95$ $120.09\!\pm\!56.34$
RADAC $764.93\!\pm\!741.86$ $1094.93\!\pm\!806.85$
HalfCheetah-m-r CQL $38.85\!\pm\!38.44$ $40.23\!\pm\!38.44$
CODAC $0.12\!\pm\!0.08$ $1.48\!\pm\!0.26$
ORAAC $315.87\!\pm\!69.27$ $161.54\!\pm\!68.76$
FlowQL $1909.57\!\pm\!395.55$ $568.43\!\pm\!256.85$
DiffusionQL $2261.16\!\pm\!531.18$ $1439.77\!\pm\!461.28$
RAFMAC $2696.61\!\pm\!110.68$ $1499.80\!\pm\!394.08$
RADAC $2674.72\!\pm\!51.76$ $1401.03\!\pm\!199.08$
Walker2d-m-r CQL $14.68\!\pm\!5.52$ $95.30\!\pm\!18.50$
CODAC $26.39\!\pm\!7.97$ $36.56\!\pm\!12.92$
ORAAC $160.23\!\pm\!147.55$ $359.49\!\pm\!302.72$
FlowQL $647.33\!\pm\!166.12$ $29.64\!\pm\!110.73$
DiffusionQL $23.50\!\pm\!4.44$ $53.55\!\pm\!12.30$
RAFMAC $778.00\!\pm\!130.03$ $7.92\!\pm\!35.77$
RADAC $383.87\!\pm\!288.95$ $309.70\!\pm\!246.62$
Hopper-m-r CQL $2.28\!\pm\!42.17$ $130.48\!\pm\!53.25$
CODAC $3.61\!\pm\!18.41$ $105.41\!\pm\!19.86$
ORAAC $30.00\!\pm\!32.77$ $179.92\!\pm\!61.46$
FlowQL $448.26\!\pm\!70.39$ $33.21\!\pm\!43.38$
DiffusionQL $22.15\!\pm\!24.93$ $163.82\!\pm\!59.18$
RAFMAC $350.36\!\pm\!33.05$ $36.69\!\pm\!28.35$
RADAC $453.64\!\pm\!68.46$ $87.04\!\pm\!123.96$
Pareto Frontier Analysis: Return vs. Safety Violations
Figure 8 plots mean return (y) against safety-violation counts per episode (x), with color indicating training progress. Unless noted, comparisons refer to the same 1000-step evaluation as in Table 4. We organize the discussion by algorithm.
Across datasets, RADAC populates the upper-left region of the frontier: for comparable return, it tends to incur fewer violations. Only for halfcheetah–medium–expert, RADAC sometimes drifts up-right (higher return with slightly more violations) because the penalty is light and non-terminating, so near-threshold speed pays off, consistent with its best Mean/CVaR. Mechanistically, diffusion with CVaR guidance enables fine-grained reweighting away from safety thresholds while BC keeps samples on-manifold, so trajectories in the plot drift left (fewer violations) without sacrificing return. RAFMAC pushes the top of the frontier in mean—most clearly on walker2d and halfcheetah-m-r—and is competitive in CVaR (Table 4). Deterministic ODE transport with low-variance policy gradients and velocity-matching BC yields fast on-manifold improvement; in the Pareto view this appears as high-return points with modest violation counts. Because the transport is geometry-preserving, boundary-adjacent mass tends to thin rather than disappear abruptly; CVaR improves as the velocity field adapts. ORAAC forms the frontier on hopper-m-e with few violations and strong returns, matching its leading scores under terminating pose hazards. In other settings it remains reliably conservative (low violations) at the cost of mean on some tasks, consistent with anchor-based regularization. FlowQL often achieves high-mean points but with comparatively higher violation counts in the Pareto plot. Without tail-aware guidance, safety depends on the expected-value critic and task smoothness, explaining the variability across datasets. DiffusionQL exhibits wider scatter: runs either reach moderate returns with elevated violations or collapse to low-return, near-zero violation regions. This variability is consistent with value-only guidance under stochastic penalties and matches its weaker CVaR. CODAC clusters in the low-return/low-violation corner across tasks, as expected from conservative critics.
[…]
Refer to captionMean Rewardhalfcheetah-m-e
Refer to captionhalfcheetah-m-r
Refer to captionwalker2d-m-e
Refer to captionwalker2d-m-r
Refer to captionhopper-m-e
Refer to captionhopper-m-r
 RADAC
 DiffusionQL
 FlowQL
 ORAAC
  CODAC
D.3 Ablation Study
We evaluate RADAC and RAFMAC with three risk distortions CVaR, Wang, and CPW under the same 1000‑step evaluation protocol used above. Across seeds, Wang generally tilts updates toward higher means and weaker tails; CPW sits between CVaR and Wang but shows higher variance across seeds. Overall, CVaR is the most reliable choice for lower‑tail control at comparable mean.
Table 5: Ablation (1000-step evaluation). RADAC/RAFMAC with CVaR, Wang, and CPW on halfcheetah-medium-replay and walker2d-medium-replay. Scores are mean $\pm$ s.e. over 3 seeds.
Method Distortion HalfCheetah–m–r Walker2d–m–r
Mean CVaR0.1 Mean CVaR0.1
RADAC CVaR $2758.5\pm 84.1$ $1759.5\pm 71.5$ $681.3\pm 409.3$ $395.1\pm 438.3$
RADAC Wang $2653.5\pm 86.5$ $310.8\pm 92.6$ $417.3\pm 397.0$ $52.1\pm 11.4$
RADAC CPW $2777.9\pm 93.7$ $1061.6\pm 731.7$ $64.3\pm 149.3$ $203.6\pm 69.8$
RAFMAC CVaR $2835.8\pm 116.3$ $1981.2\pm 405.3$ $698.8\pm 215.5$ $5.6\pm 60.8$
RAFMAC Wang $2625.6\pm 113.8$ $462.5\pm 427.6$ $552.2\pm 134.8$ $706.4\pm 687.5$
RAFMAC CPW $2539.2\pm 31.1$ $95.9\pm 92.3$ $360.7\pm 49.6$ $71.6\pm 22.1$
Appendix E Experimental Details
E.1 2D Synthetic Task Details
Risky–Bandit dataset
We generate $N=10^{4}$ state–action–reward tuples with dummy zero states. Actions come from two modes: (i) Ring (80%): radius $0.9\pm 0.04$; base reward $\mathcal{N}(9,\,0.3^{2})$; with probability $0.05$ a trap penalty 40-40 is applied (heavy lower tail). (ii) Centre (20%): $\mathcal{N}(\mathbf{0},\,0.1^{2}\mathbf{I})$; reward $\mathcal{N}(5,\,0.3^{2})$. Actions are clipped to $1,1]^{2}$.
All methods train on the same static dataset; when a BC regulariser is required we use the standard loss of the underlying generator. RADAC adds the CVaR term from Eq. 11 to the diffusion/flow BC objective and backpropagates. For each trained policy we draw 1,0001{,}000 action samples for visualisation in Fig. 3.
E.2 Stochastic-D4RL MuJoCo Suite
Datasets We adopt the stochastic MuJoCo protocol for risk-sensitive offline RL, following (Urpí et al., 2021). Policies are evaluated on
$\{\textsc{hopper},\ \textsc{walker2d},\ \textsc{halfcheetah}\}\ \times\ \{\textsc{medium-expert},\ \textsc{medium-replay}\},$
Compared to prior work, we prefer medium-expert and medium-replay to validate both risk sensitivity and policy expressiveness under multimodal action distributions. For training, we relabel per-transition rewards in the offline datasets to inject stochastic hazards (velocity or torso-pitch thresholds with Bernoulli penalties and early termination); the same hazard model is used at evaluation. This ensures the critic and the policy are trained on the risk-aware rewards rather than only being tested under hazards.
Settings Each task defines a monitored signal and an additive Bernoulli penalty when a safety condition is violated; pose-based tasks also include an early-termination threshold.
  • HalfCheetah : monitor forward velocity. Apply a penalty with probability $p=0.05$ if the threshold is exceeded. Thresholds/penalties: medium-expert/medium-replay uses v>10.0v>10.0/ $v>5.0$ with penalty 70.0-70.0. No early termination. Max episode steps: $200$.
[…]
    • Hopper: healthy range [0.1,0.1][-0.1,0.1] rad; penalty 50.0-50.0 when $|\tilde{\theta}|>0.1$; early termination if $|\theta|>0.2$.
    • Walker2D: healthy range [0.5,0.5][-0.5,0.5] rad; penalty 30.0-30.0 when $|\tilde{\theta}|>0.5$; early termination if $|\theta|>1.0$.
E.3 Baselines: Implementation & Hyperparameters
We include five representative offline-RL methods standard:
  • CODAC (Ma et al., 2021) (distributional conservative learning). We primarily use the CVaR-optimizing specification (“CODAC-C”, $\mathrm{CVaR}_{0.1}$ objective).
  • ORAAC (Urpí et al., 2021) (offline risk-averse actor–critic). A distributional critic with imitation-regularized policy optimizing a coherent risk objective.
  • CQL (Kumar et al., 2020) (value pessimism). Non-distributional conservative Q-learning baseline.
  • DiffusionQL (Wang et al., 2023) (expressive risk-neutral diffusion policy).
  • FlowQL (Park et al., 2025) (expressive risk-neutral flow-matching policy).
Hyperparameter selection & tuning
For each of baselines, we run all baselines ourselves and tune the following parameters or adopt authors’ recommended settings, mirroring the practice in Ma et al. (2021); Urpí et al. (2021); Wang et al. (2023); Park et al. (2025); Kumar et al. (2020).
  • FlowQL (Park et al., 2025): we sweep the policy weight $\alpha\in\{1,10,30,100,1000\}$ per task and report the best-performing setting (selection by $\mathrm{CVaR}_{0.1}$ unless noted).
  • DiffusionQL (Wang et al., 2023): we consider $\eta\in\{0.1,0.5,1.0\}$ for BC coefficient . we use authors’ recommended configuration for other parameters without retuning. We also used the best checkpoint of their model on each benchmark by following their protocol.
  • ORAAC (Urpí et al., 2021): use the paper’s recommended configuration (distributional critic, risk level $\alpha=0.1$, anchor/prior regularization) without additional sweeps.
  • CODAC (Ma et al., 2021): use the paper’s tuned settings for D4RL (risk level $\alpha=0.1$) without further tuning.
  • CQL (Kumar et al., 2020): use the standard conservative coefficient and implementation defaults for MuJoCo locomotion.
E.4 Estimating the OOD Visitation Rate with a 1 –NN Detector
At evaluation time, we estimate the fraction of actions produced by a policy that fall outside the empirical action support of the offline dataset. This fraction is reported as the OOD action rate $\varepsilon_{\text{act}}$ in Sec. 5.3. Let $\mathcal{A}_{\mathcal{D}}=\{a_{i}\}_{i=1}^{N}$ be the set of actions in the offline dataset (per task), and let $\|\cdot\|_{2}$ denote the Euclidean norm in action space. In our MuJoCo tasks, actions are already scaled to [1,1][-1,1] per dimension; we therefore use $\ell_{2}$ directly.
For each dataset action $a_{i}$, compute its nearest neighbour among the other dataset actions and the associated distance
$d_{i}\;=\;\min_{j\neq i}\,\|a_{i}-a_{j}\|_{2}.$
Define the robust scale
$\operatorname{medNN}\;=\;\mathrm{median}\{d_{i}\}_{i=1}^{N}.$
We set the OOD threshold to $\tau\,=\,\kappa\cdot\operatorname{medNN}$ with $\kappa{=}3$ unless otherwise stated.
Label evaluation actions
Let $\mathcal{A}_{\text{eval}}=\{a^{(\text{eval})}_{t}\}_{t=1}^{T}$ be all actions emitted across evaluation rollouts (we use 5 seeds $\times$ 10 episodes/seed by default). For each $a^{(\text{eval})}_{t}$, compute its distance to the dataset action set
$d^{(\text{eval})}_{t}\;=\;\min_{i\in\{1,\dots,N\}}\|a^{(\text{eval})}_{t}-a_{i}\|_{2},$
and assign the indicator
$\mathbf{1}_{\text{OOD}}(a^{(\text{eval})}_{t})\;=\;\mathbb{I}\!\left\{d^{(\text{eval})}_{t}\;>\;\tau\right\}.$
In practice we compute $d^{(\text{eval})}_{t}$ via a KD–tree built on $\mathcal{A}_{\mathcal{D}}$ (query time $O(\log N)$ in low to moderate dimensions).
We define the OOD action rate as the per–action frequency
$\varepsilon_{\text{act}}\;=\;\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}_{\text{OOD}}(a^{(\text{eval})}_{t})$.
We report the mean and standard error over 5 seeds. Because episodes may terminate early under the stochastic wrappers, $T$ is the actual number of executed timesteps (not a fixed horizon), which makes the rate comparable across seeds.
Confidence intervals
For a seed–level rate $\hat{\varepsilon}$ with $T$ trials, we use a binomial approximation for the standard error $\mathrm{SE}=\sqrt{\hat{\varepsilon}(1-\hat{\varepsilon})/T}\,$ and report the across–seed mean $\pm$ s.e.
Table 6: RAMAC: hyperparameters. We keep only the knobs that materially affect performance and stability. Values are our defaults; brackets show typical sweep ranges.
Global
Discount $\gamma$ $0.99$
Batch size $B$ $256$
Target update $\tau_{\text{target}}$ $0.005$
Risk level $\alpha$ $0.1$
Critic (Deterministic IQN)
#Quantiles NN $32$
Grid $\mathcal{T}_{N}$ $\{(i-\tfrac{1}{2})/N\}_{i=1}^{N}$ (fixed)
Embedding dim $128$
Critic LR $3\times 10^{-4}$
Huber $\kappa$ 11 (fixed)
Double IQN enabled
Actor (shared)
Actor LR $3\times 10^{-4}$
BC weight $\lambda_{\mathrm{BC}}$ $1.0$
Risk weight $\eta$ RADAC: $0.05$   [0.020.020.10.1],  RAFMAC: $1000$   [10010010001000]
Double critic clipping RADAC:[150150150150]-[$300300$],  RAFMAC:[$300300$]
RADAC-specific
Reverse diffusion steps $T$ 55 (VP schedule)
RAFMAC-specific
Flow steps $K$ $10$ (Euler, $\Delta t{=}1/K$)